OpenStudy (anonymous):

for the question -4y=8x, what is the constant of variation? a. -4 b. -2 c. 1 d. 2 Suppose y varies directly with x, and y=24 when x=8. what is the value of y when x=10? a. 10 b. 30 c. 80 d. 240 @Ahaanomegas

OpenStudy (anonymous):

For the first one, what do you define as the constant of variation? y/x? If so, divide out to get -2. For #2, Note that \[ \frac {y}{x} \] is a constant, and this constant is 24/8 = 3, so the answer is 10*3 = 30.

OpenStudy (mathstudent55):

Solve the first one for y.

OpenStudy (mathstudent55):

The constant of variation is the slope. Solve the equation for y to get it in the form y = mx + b. In your case, b = 0. m is the slope.